The steering wheel is turned and held in place so that the car travels in a perfect circle in the counterclockwise direction. The car travels in a complete circle and stops. The front wheels are 2 m apart, and the area between the two circular tracks made by the two front wheels is 200 square metres. What is the area of the circle traced by the left front wheel in square metres? (Assume that pi=3.14)






choose the answer and explain the reason to choose your response
698.98
847.02
3083.88
2986.23


Answer by stanbon(75887)

The steering wheel is turned and held in place so that the car travels in a perfect circle in the counterclockwise direction. The car travels in a complete circle and stops. The front wheels are 2 m apart, and the area between the two circular tracks made by the two front wheels is 200 square metres. What is the area of the circle traced by the left front wheel in square metres? (Assume that pi=3.14)
------
Draw the figure of one circle inside another.
Inner circle area = (pi)r^2
Outer circle area = (pi)(r+2)^2
---
Equation:
Outer Area - Inner area = 200 m^2
(pi)[(r+2)^2-r^2] = 200
4r+4 = 200/3.14
r = 14.92 meters
----
Inner circle Area = (pi)14.92^2 = 699.31 sq. meters
=====================================================
Cheers,
Stan H.





choose the answer and explain the reason to choose your response
698.98
847.02
3083.88
2986.23